OpenStudy (anonymous):

Point A is located at (-2,2) and D is located at (4,-2) find the coordinates of the point that lies halfway between A and D

OpenStudy (anonymous):

\[\left( \frac{ x _{1}+x _{2} }{ 2 },\frac{ y _{1}+y _{2} }{ 2 }\right)\] Mid point